OpenStudy (anonymous):

How many seven-letter words contain at least two X's? Hint: The Bad ones are those with no X's and those with exactly one X. Think carefully about counting the number of words with exactly one X.

OpenStudy (ash2326):

@tim93 Find the no. of words with no X's?

OpenStudy (anonymous):

= all possible combinations = all possible combinations with no X = all possible combinations with 1 X Any thoughts for any of them ?

OpenStudy (ash2326):

@tim93 yeah try this after you get up:D Night:)

OpenStudy (anonymous):

All possible combinations = 26^7 All possible combinations with no x = 25^7 All possible combinations with only one x = ?????? getting stuck here

OpenStudy (anonymous):

word means any string of 7 letters?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no \(x\) them would be \(25^7\) i guess

OpenStudy (anonymous):

maybe only one x would be 26^6 + 26^7??

OpenStudy (anonymous):

one \(x\) would be \(7\times 25^6\) i think here is my reasoning, see if it make sense x in the first spot, leaving 6 places and 25 choices or x in the 2nd spot, leaving 6 places and 25 choices or ...

OpenStudy (anonymous):

wait,

OpenStudy (anonymous):

26 * 6^25 for wods with only one x?

OpenStudy (anonymous):

With only one x would be 26^6 to choose the other 6 letters, and then multiply that by 7 because the x could be placed 7 different spots.

OpenStudy (anonymous):

you have the base and exponent backwards i think

OpenStudy (anonymous):

so the final equation would be ... 26^7 - 25^7 - (26 * 25^6) ??? sound right?

OpenStudy (anonymous):

x, __, __, __, __, __, __ 6 slots in which each has 25 choices

OpenStudy (anonymous):

ok \(26^7\) is the total with no restriction right ?

OpenStudy (anonymous):

with the restriction that there is no x it is \(25^7\)

OpenStudy (anonymous):

and with the restriction that there is one x i still believe by what i wrote above that it is \(7\times 25^7\)

OpenStudy (anonymous):

maybe i am wrong, track record not so good today

OpenStudy (anonymous):

26^7 - 25^7 - (7* 25^7)

OpenStudy (anonymous):

oh yes i am wrong \[7\times 25^6\]

OpenStudy (anonymous):

Tim, to create a word with only 1 x, you must choose the other 6 letters. For each of those, you have 25 options, so to choose those letters, you get this many possibilities 25^6 Now, you've chosen all of the other 6 letters, but you can still get different choices by placing the x in different spots. For example, xaaaaaa axaaaaa aaxaaaa So, that means there are 7 different spots to place the x, therefore the total number of possibilities is 7*25^6

OpenStudy (zarkon):

\[\sum_{x=2}^{7}{7\choose x}25^{7-x}\] \[=26^7-\sum_{x=0}^{1}{7\choose x}25^{7-x}\]

OpenStudy (anonymous):

zarkon gives direct computation

OpenStudy (anonymous):

26^7 - 25^7 - (7 * 25^6)

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

what smoothmath said

OpenStudy (anonymous):

thanks guys